# Generative Physical AI in Vision: A Survey

### v1

@@ 4 Physical Understanding
Physical understanding in machine learning aims to infer underlying physical models, laws, or parameters from observational data such as images and videos. This capability is crucial for tasks that require modeling real-world dynamics, including object motion, material behavior, and environmental interactions. While extensive research has been dedicated to physical reasoning and model discovery[126, 127, 128, 129, 130, 131, 132, 133, 134, 135, 136, 137, 138, 139, 140, 141, 142, 143, 144, 145, 146, 147, 148, 149, 150, 151, 152, 153, 154, 155], this area is largely orthogonal to physics-aware generative modeling, which is thus out of the core focus of this survey. Nonetheless, accurately estimating physical parameters remains a foundational component in enabling physically consistent generation. In the context of physics-aware generative AI, determining the physical parameters that drive simulations is essential for producing realistic and physically plausible outputs. We identify three major approaches to obtain physical parameters as below.
@@ 5 Physics-Aware Generation
In this section, we first provide an overview of the primary classes of generative models employed in physics-aware generation. We then introduce the main paradigms for incorporating physics-based constraints and knowledge into these generative models.
@@ 5.1 Generative Models
We begin by introducing widely used models including Generative Adversarial Networks (GANs), Diffusion Models (DMs), and neural rendering methods—including Neural Radiance Field (NeRF) and Gaussian Splatting (GS)—which, although traditionally viewed as rendering or reconstruction frameworks, are interpreted as generative models in a broader sense in this survey. These generative models can handle a variety of data modalities (e.g., images, videos, 3D, 4D, or interactive environments) and encompass various domains (e.g., general data, human-centric data, and indoor scenes).
@@ 5.2 Generative Adversarial Networks
A generative adversarial network[3] consists of two neural networks—a generator and a discriminator—engaged in a competitive process. The generator seeks to create synthetic data that closely resembles real-world data, while the discriminator attempts to distinguish between real and generated data. This adversarial dynamic drives both networks to improve iteratively, resulting in a generator capable of producing highly realistic data samples. Specifically, the generator $G$ and the discriminator $D$ play a two-player minimax game with the following loss function:
@@ 5.3 Diffusion Models
where $\sigma$ is the variance schedule. This denoising process can be parameterized in several ways, including mean prediction, noise prediction, and clean data prediction[76]. This iterative refinement allows diffusion models to generate detailed and diverse outputs while maintaining strong theoretical foundations. Score-based models are of an equivalent formulation as diffusion models but from the differential equation perspective[160, 161, 162]. The denoising process can be accelerated by samplers such as DDIM[4] and DPM-Solvers[163]. Latent diffusion models[164] conduct the forward and reverse processes in a high-level latent space and delegate low-level reconstruction to an auxiliary VAE. Techniques like classifier guidance and classifier-free guidance further improve controllability and generation quality[165]. Diffusion model and its later development[166, 167, 168, 169, 170, 171, 172] have been greatly advanced generative tasks for visual data.
@@ 5.4 Neural Radiance Field
By optimizing the network to minimize the error in rendering the observed views of a scene, NeRF can synthesize photorealistic novel views by querying the MLP along camera rays and using volume rendering techniques to project the output colors and densities into an image. PixelNeRF[173] enables neural scene representation from only one single observed image. MIP-NeRF[174] reduces aliasing artifacts and significantly improves the ability to represent fine details. DNeRF[175] extends neural radiance fields to dynamic objects with motions. These advancements have propelled the neural radiance field into widespread applications in physics-informed scene representations for dynamic and large-scale environments.
@@ 5.6 Physics-Aware Generation w/ Explicit Simulation
In the following, we review papers working on physics-aware generation with explicit physical simulation (PAG-E), which are grouped into six categories according to the paradigm of how physics simulation (Sim) is integrated with the generative model (Gen). Note that one paper can usually couple multiple paradigms, but is classified based on the most evident one.
@@ 5.6.1 Paradigm 1: Gen-to-Sim (GtS)
This category of papers usually appends physical properties to a generative representation in a post-processing manner to make it simulatable and interactable.
Simulation Elements in NeRF Field. PIE-NeRF[100] uses Poisson disk sampling to distribute “particles" in the NeRF density field, which form simulation elements through Voronoi grouping. Quadratic generalized moving least square (Q-GMLS) strategy and Lagrangian dynamics are then applied to the elements for simulation, enabling users to interact with the NeRF scene using external forces. Video2Game[90] automatically converts a single video of a real-world scene into an interactive virtual environment. To achieve this, a NeRF-generated scene is segmented into individual objects. Each segmented object is then assigned physical parameters such as mass, friction, and collision geometry (e.g., sphere, box, convex polygon) to be simulated with rigid-body physics in real time using WebGL-based game engine.
Gaussian Blobs as Simulation Elements. PhysGaussian[96] first builds a Gaussian Splatting with anisotropic regularization, and then imparts physically grounded behavior to the Gaussian kernels through Material Point Method (MPM). The method applies continuum mechanics to evolve 3D Gaussian kernels, where the kernels are treated as discrete particle clouds for discretizing the continuum. This allows the system to model deformation realistically by tracking physical quantities like stress and strain. Similarly, GASP[103] also leverages the Material Point Method (MPM) to simulate physical behaviors. GASP converts Gaussian components into triangle-based meshes, on which the MPM is applied, and the results are re-parameterized back into Gaussian components for rendering. Spring-Gau[99] first reconstructs static Gaussian, from which anchor points are sampled and are connected via “springs" to model elastic behavior. The model uses differentiable simulation to optimize learnable stiffness and damping parameters from video observations, to transform the Gaussian representation to be dynamic and simulatable. Phy124[112] generates physics-consistent 4D content from a single image, by evolving dynamic 3D content adhering to natural physical laws. Specifically, 3D Gaussians are generated from a single image using diffusion prior and then animated by attaching an MPM simulator.
@@ 5.6.2 Paradigm 2: Sim-in-Gen (SiG)
Conditions from Simulation. In GPT4Motion[89], the Blender-based physics simulation is integrated into the video generation pipeline. Specifically, GPT-4 translates user prompts into Python scripts that control Blender’s physics engine to simulate realistic motion. The output of Blender—rendered as edge and depth maps—is then fed into ControlNet-modified Stable Diffusion to generate videos that are both visually consistent and physically plausible. MotionCraft[95] enhances pretrained image diffusion models and introduces motions by warping the noise latent space with optical flow derived from physics simulation. This process allows the model to generate temporally consistent frames that evolve according to physically accurate dynamics such as fluid dynamics, rigid body motion, and multi-agent interactions. Similarly, by using image-based warping with simulated motion dynamics, PhysGen[87] generates physically plausible video outputs given an input image and user-defined forces and torques. The simulation is based on rigid-body dynamics governed by Newton’s Laws, using physical parameters inferred by large foundation models.
@@ 5.6.4 Paradigm 4: Sim-Constrained Gen (ScG)
Simulation for Data Filtering DiffuseBot[92] employs simulation as a component for data filtering during its embedding optimization phase. Specifically, 3D robot models generated from diffusion models are evaluated using a differentiable physics simulation to assess the performance of each robot based on task-specific metrics. A filtering mechanism then selects the top-performing designs to be retained, gradually skewing the sampling distribution of the generative model toward successful robot designs.
Reinforcement Learning with Physical Feedback. Furuta et al.[157] addresses the challenge of generating realistic dynamic object interactions in text-to-video models, by using reinforcement learning (RL) fine-tuning from external feedback—particularly from vision-language models (VLMs). The VLM-based feedback mechanism serves as a form of simulation or evaluation of physical realism, which constrains and guides the generation model toward producing physically plausible videos.
@@ 5.6.5 Paradigm 5: Gen-Constrained Sim (GcS)
Score Distillation Sampling. Physics3D[97] combines elastoplastic and viscoelastic MPM with a video diffusion model. To optimize physical parameters, it utilizes Score Distillation Sampling (SDS) from a pretrained video diffusion model. The diffusion model provides motion pattern priors learned from video data, guiding the simulation toward physically plausible behaviors without requiring manual specification of material properties. DreamPhysics[98] further proposes Motion Distillation Sampling (MDS) to better capture motion-specific priors and reduce color bias in the optimization.
Learning Physical Parameters from Generated Data. PhysDreamer[101] synthesizes realistic interactive dynamics for objects by extracting physical dynamics priors from video generation models. Different from score distillation sampling, this method learns such priors from generated videos. The method first uses a pre-trained image-to-video generation model to create a reference video of a static object undergoing plausible motions. Then it optimizes physical parameters such as Young’s modulus in MPM simulation by maximizing the visual similarity between the simulation-rendered video and the generated reference video.
@@ 5.7 Physics-Aware Generation w/o Explicit Simulation
In this section, we introduce representative works in generative AI for vision that exhibit a degree of physical awareness without relying explicitly on physical simulation (PAG-I).
@@
Physical Awareness in Large Video Models. Recent advancements in large-scale video generation models, such as OpenAI’s Sora[20], OpenSora[179], CogVideoX[180], VideoCrafter[29], LAVIE[35], ModelScope[181], demonstrate emergent physical reasoning capabilities. These models, scaled up and trained on vast datasets of internet-scale video content, have shown promising abilities to implicitly capture and replicate certain physical dynamics and causal relationships present in the real world[182]. This emergent physical understanding enables them to generate complex scenes with coherent object interactions, realistic motions, and plausible temporal dynamics. This progress positions them as potential foundations for universal simulators capable of modeling diverse physical phenomena, despite lacking explicitly programmed physics engines[45]. However, this implicit physical reasoning remains in its early stage and systematic evaluations are still open areas. For example, PhyGenBench[183] introduces a comprehensive benchmark designed to evaluate the physical commonsense understanding of text-to-video generation models. The findings highlight that current models, despite their impressive visual generation capabilities, struggle to represent basic physical laws accurately. Kang et al.[184] investigate whether large-scale video generation models can autonomously learn and generalize physical laws solely from visual data. It is found that scaling model size and dataset do not help with out-of-distribution physical generalization. Current models primarily rely on the existence of closely similar training examples rather than abstracting general physical rules, underscoring the need for more targeted methods to develop fully physics-aware generative models.
@@
Physical Awareness from Large Language Models. Large language models can also supply physical knowledge for improving the generation of visual data. PhyT2V[185] addresses the lack of physical realism in text-to-video models by integrating large language models to refine textual prompts iteratively. The process first extracts objects and relevant physical rules from the initial text prompt, and then uses a video captioning model to generate semantic descriptions of the generated video and compare it with the original prompt to identify mismatches. Finally, the text prompt is refined based on the extracted physical properties to correct the mismatches.
@@
Generative Interactive Dynamics. Generative interactive dynamics focus on modeling the natural dynamics of objects without requiring explicit physical simulations, implicitly learning the physical relationships and interactions. Blattmann et al.[186] makes an early attempt to predict object dynamics in response to localized user interactions within an image-to-video formulation. Generative Image Dynamics[187] generates realistic scene motion from a single image, via a latent diffusion model that predicts spectral volumes representing dense long-term pixel trajectories. Motion-controlled video generation and video editing belong to another form of generative dynamics[188, 57]. Motion Prompting[56] leverages motion trajectories as conditioning signals, allowing for fine-grained and flexible control of local and global scene motions. Similarly, VideoComposer[43] and Yoda[55] utilize motion vectors as explicit temporal control signals to allow for precise capture of inter-frame dynamics and patterns. Motion Dreamer[189] decouples motion reasoning from high-fidelity video generation, using intermediate motion representations such as optical flow, instance segmentation maps, and depth maps. Motion Guidance[190] controls the diffusion generation process using gradients derived from optical flow estimators, enabling detailed motion edits without model retraining.

### v2

@@ 4 Physical Understanding
Physical understanding in machine learning is the process of inferring the underlying physical models, laws, or parameters from observational data such as images and videos. This capability is crucial for tasks that require modeling real-world dynamics, including object motion, material behavior, and environmental interactions. While extensive research has been dedicated to physical reasoning and model discovery[86, 87, 88, 89, 90, 91, 92, 93, 94, 95, 96, 97, 98, 99, 100, 101, 102, 103, 104, 105, 106, 107, 108, 109, 110, 111, 112, 113, 114, 115], this area is largely orthogonal to physics-aware generative modeling, and hence outside the core focus of this survey. Nonetheless, accurately estimating physical parameters remains a foundational component in enabling physically consistent generation. In the context of physics-aware generative learning, determining the physical parameters that drive simulations is essential for producing realistic and physically plausible outputs. We identify three major approaches to obtain physical parameters as below.
@@ 5 Physics-Aware Generation
We first provide an overview of the primary classes of generative models employed in physics-aware generation and then introduce the main paradigms for incorporating physics-based constraints and knowledge into these generative models.
@@ 5.1 Generative Models
We begin by introducing widely used models including Generative Adversarial Networks (GANs), Diffusion Models (DMs), and neural rendering methods including Neural Radiance Field (NeRF) and Gaussian Splatting (GS) which although traditionally viewed as rendering or reconstruction frameworks, are interpreted as generative models in a broader sense in this survey. These generative models can handle a variety of data modalities (e.g., images, videos, 3D, 4D, or interactive environments) and encompass various domains (e.g., general data, human-centric data, and indoor scenes).
@@ 5.2 Generative Adversarial Networks
A Generative Adversarial Network[3] consists of two neural networks, i.e., a generator and a discriminator, engaged in a competitive process. The generator seeks to create synthetic data that closely resembles real-world data, while the discriminator attempts to distinguish between real and generated data. This adversarial dynamic drives both networks to improve iteratively, resulting in a generator capable of producing highly realistic data samples. Specifically, the generator $G$ and the discriminator $D$ play a two-player minimax game with the following loss function:
@@ 5.3 Diffusion Models
where $\sigma$ is the variance schedule. This denoising process can be parameterized in several ways, including mean prediction, noise prediction, and clean data prediction[78]. This iterative refinement allows diffusion models to generate detailed and diverse outputs while maintaining strong theoretical foundations. Score-based models are of an equivalent formulation as diffusion models but from the differential equation perspective[155, 156, 157]. The denoising process can be accelerated by samplers such as DDIM[4] and DPM-Solvers[158]. Latent diffusion models[159] conduct the forward and reverse processes in a high-level latent space and delegate low-level reconstruction to an auxiliary VAE. Techniques like classifier guidance and classifier-free guidance further improve controllability and generation quality[160]. Diffusion model and its later developments[161, 162, 163, 164, 165, 166, 167] have greatly advanced generative tasks for visual data.
@@ 5.4 Neural Radiance Field
By optimizing the network to minimize the error in rendering the observed views of a scene, NeRF can synthesize photorealistic novel views by querying the MLP along camera rays and using volume rendering techniques to project the output colors and densities into an image. PixelNeRF[168] enables neural scene representation from only one single observed image. MIP-NeRF[169] reduces aliasing artifacts and significantly improves the ability to represent fine details. DNeRF[170] extends neural radiance fields to dynamic objects with motions. These advancements have propelled the neural radiance field into widespread applications in physics-informed scene representations.
@@ 5.6 Physics-Aware Generation w/ Explicit Simulation
In the following, we review papers working on physics-aware generation with explicit physical simulation (PAG-E), which are grouped into six categories (Fig. 5) according to the paradigm of how physics simulation (Sim) is integrated with the generative model (Gen). Note that one paper may follow multiple paradigms, and we assign it to the most relevant one in such cases.
Y1.0e0]ptPhysics-Aware Generation with Explicit Simulation Paradigm 1 (Sec.5.6.1) Gen-to-Sim Applications in VR and Robotics VR-GS[118], LIVE-GS[142], DreMa[119] Physical Feature Field Feature Splatting[141], Phys4DGen[140], SimAnything[139] Gaussian Blobs as Simulation Elements PhysGaussian[117], GASP[120], Spring-Gau[133], Phy124[121], DecoupledGS [150] Simulation Elements in NeRF PIE-NeRF[116], Video2Game[143] Paradigm 2 (Sec.5.6.2) Sim-in-Gen Generative Programs for Simulation AutoVFX[148], GPT4Motion[145] Simulation as Optimization PhyCAGE[124], PhysDiff[138] Conditions from Simulation GPT4Motion[145], MotionCraft[127], PhysGen[146], PhysAnimator [152] Paradigm 3 (Sec.5.6.3) Gen-and-Sim Alternating Simulation and Generation PhysMotion[147] Joint Geometry and Physics Learning PAC-NeRF[131], iPAC-NeRF[132] Paradigm 4 (Sec.5.6.4) Sim-Constrained Gen Training Reference Model using Simulation Data Zhao et al. [21] Post-training with Physical Feedback Furuta et al.[144], DSO [151] Simulation for Data Filtering DiffuseBot[137] Loss Functions from Simulation PhysComp[123], PhyRecon[125], Atlas3D[126], Mezghanni et al.[122] Paradigm 5 (Sec.5.6.5) Gen-Constrained Sim Learning Physical Parameters from Generated Data PhysDreamer[134] Score Distillation Sampling Physics3D[136], DreamPhysics[135], PhysFlow [149] Paradigm 6 (Sec.5.6.6) Sim-evaluated Gen Liu et al.[128], PhysPart[129], PhyScene[130]
Figure 5: Structure of Paradigms, Ideas, and Methods in Physics-Aware Generation with Explicit Physical Simulation.
@@ 5.6.1 Paradigm 1: Gen-to-Sim (GtS)
This category of methods usually appends physical properties to a generative representation in a post-processing manner to make it simulatable and interactable.
Simulation Elements in NeRF. PIE-NeRF[116] uses Poisson disk sampling to distribute “particles" in the NeRF density field, which form simulation elements through Voronoi grouping. Quadratic generalized moving least square (Q-GMLS) strategy and Lagrangian dynamics are then applied to the elements for simulation, enabling users to interact with the NeRF scene using external forces. Video2Game[143] automatically converts a single video of a real-world scene into an interactive virtual environment. To achieve this, a NeRF-generated scene is segmented into individual objects. Each segmented object is then assigned physical parameters such as mass, friction, and collision geometry (e.g., sphere, box, convex polygon) to be simulated with rigid-body physics in real time using WebGL-based game engine.
Gaussian Blobs as Simulation Elements. PhysGaussian[117] first builds a Gaussian Splatting with anisotropic regularization, and then imparts physically grounded behavior to the Gaussian kernels through Material Point Method (MPM). The method applies continuum mechanics to evolve 3D Gaussian kernels, where the kernels are treated as discrete particle clouds for discretizing the continuum. This allows the system to model the deformation realistically by tracking physical quantities like stress and strain. Similarly, GASP[120] also leverages the Material Point Method (MPM) to simulate physical behaviors. GASP converts Gaussian components into triangle-based meshes, on which the MPM is applied, and the results are re-parameterized back into Gaussian components for rendering. Spring-Gau[133] first reconstructs static Gaussian, from which anchor points are sampled and are connected via “springs" to model elastic behavior. The method uses differentiable simulation to optimize learnable stiffness and damping parameters from video observations, to transform the Gaussian representation to be dynamic and simulatable. Phy124[121] generates physics-consistent 4D content from a single image, by evolving dynamic 3D content adhering to natural physical laws. Specifically, 3D Gaussians are generated from a single image using diffusion prior and then animated by attaching an MPM simulator. DecoupledGS [150] separates objects and contact surfaces during scene restoration to improve the realism of interactive physical simulations such as user-specified external forces, scene collisions, and object fracturing.
@@ 5.6.2 Paradigm 2: Sim-in-Gen (SiG)
Conditions from Simulation. In GPT4Motion[145], Blender-based physics simulation is integrated into the video generation pipeline. Specifically, GPT-4 translates user prompts into Python scripts that control Blender’s physics engine to simulate realistic motion. The output of Blender, rendered as edge and depth maps, is then fed into ControlNet-modified Stable Diffusion to generate videos that are both visually consistent and physically plausible. MotionCraft[127] enhances pretrained image diffusion models and introduces motions by warping the noise latent space with optical flow derived from physics simulation. This process allows the model to generate temporally consistent frames that evolve according to physically accurate dynamics such as fluid dynamics, rigid body motion, and multi-agent interactions. Similarly, by using image-based warping with simulated motion dynamics, PhysGen[146] generates physically plausible video outputs given an input image and user-defined forces and torques. The simulation is based on rigid-body dynamics governed by Newton’s Laws, using physical parameters inferred by large foundation models. PhysAnimator [152] further extends to deformable body dynamics by using mesh reconstruction and simulation to warp a sketch image. The warped sketches are then sent into ControlNet to guide the video diffusion model for generation.
@@ 5.6.4 Paradigm 4: Sim-Constrained Gen (ScG)
Simulation for Data Filtering. DiffuseBot[137] employs simulation as a component for data filtering during its embedding optimization phase. Specifically, 3D robot models generated from diffusion models are evaluated using a differentiable physics simulation to assess the performance of each robot based on task-specific metrics. A filtering mechanism then selects the top-performing designs to be retained, gradually skewing the sampling distribution of the generative model toward successful robot designs.
Post-training with Physical Feedback. Furuta et al.[144] addresses the challenge of generating realistic dynamic object interactions in text-to-video models, by using reinforcement learning (RL) fine-tuning from external feedback, particularly from vision-language models (VLMs). The VLM-based feedback mechanism serves as a form of simulation or evaluation of physical realism, which constrains and guides the generation model toward producing physically plausible videos. DSO [151] creates 3D assets using a base image-to-3D generation model and annotates their self-standing stability through physics-based simulation. The base model is then fine-tuned on the created dataset using either Direct Preference Optimization (DPO) or Direct Reward Optimization (DRO) to encourage stable samples.
Training Reference Model using Simulation Data. Zhao et al. [21] improve video generation by introducing a reference model, which is an extra video generation model in parallel with the primary generation model, but trained on synthetic data from simulation engines. The authors propose a strategy called SimDrop, where the auxiliary reference model guides the primary model to minimize artifacts unique to synthetic data while retaining the desired physical fidelity. To achieve this, a weighted combination of the primary model’s output and the reference model’s output is used in a manner similar to the classifier-free guidance of diffusion models during inference.
@@ 5.6.5 Paradigm 5: Gen-Constrained Sim (GcS)
Score Distillation Sampling. Physics3D[136] combines elastoplastic and viscoelastic MPM with a video diffusion model. To optimize physical parameters, it utilizes Score Distillation Sampling (SDS) from a pretrained video diffusion model. The diffusion model provides priors learned from video data, guiding the simulation toward physically plausible behaviors without requiring manual specification of material properties. DreamPhysics[135] further proposes Motion Distillation Sampling (MDS) to better capture motion-specific priors and reduce color bias in the optimization. Similarly, PhysFlow [149] utilizes optical flow loss to distil the knowledge from video diffusion models to optimize the material properties intialized by multimodal foundation models.
Learning Physical Parameters from Generated Data. PhysDreamer[134] synthesizes realistic interactive dynamics for objects by extracting the priors of physical dynamics from video generation models. Different from score distillation sampling, this method learns such priors from generated videos. The method first uses a pre-trained image-to-video generation model to create a reference video of a static object undergoing plausible motions. Then it optimizes physical parameters such as Young’s modulus in MPM simulation by maximizing the visual similarity between the simulation-rendered video and the generated reference video.
@@ 5.6.6 Paradigm 6: Sim-evaluated Gen (SeG)
Physics-Aware Generation w/o Explicit Simulation Physical Awareness Emergent in Large Video Models Sora [22], OpenSora [174], CogVideoX [175], ModelScope [176], Cosmos [26], etc. Physical Awareness from Large Language Models PhyT2V[177], VideoAgent [178], etc. Physical Awareness from Physics-rich Training Data WISA [20], PISA [179], etc. Generative Interactive Dynamics and Motion Controls Blattmann et al.[180], Generative Image Dynamics[181], Motion Prompting[58], VideoComposer[46], Yoda[57], Motion Dreamer[182], Motion Guidance[183], LivePhoto[184], etc. Physical Domain Data Generation CoCoGen [185], Cao et al.[186], etc.
Figure 6: Structure of Ideas and Methods in Physics-Aware Generation without Explicit Physical Simulation.
@@ 5.7 Physics-Aware Generation w/o Explicit Simulation
In this section, we introduce representative works of generative learning in vision that exhibit a degree of physical awareness without relying explicitly on physical simulation (PAG-I).
@@
Physical Awareness Emergent in Large Video Models. Recent advancements in large-scale video generation models, such as Sora [22], OpenSora [174], CogVideoX [175], VideoCrafter [32], LAVIE [38], ModelScope [176], and Cosmos [26] demonstrate emergent physical reasoning capabilities. These models, scaled up and trained on vast datasets of internet-scale video content, have shown promising abilities to implicitly capture and replicate certain physical dynamics and causal relationships present in the real world [187]. This emergent physical understanding enables them to generate complex scenes with coherent object interactions, realistic motions, and plausible temporal dynamics. This progress positions them as potential foundations for universal simulators capable of modeling diverse physical phenomena, despite the lack of of explicitly programmed physics engines[188]. However, this implicit physical reasoning remains in its early stage and systematic evaluations are still open areas. For example, PhyGenBench[189] introduces a comprehensive benchmark designed to evaluate the physical common sense understanding of text-to-video generation models. The findings highlight that current models, despite their impressive visual generation capabilities, struggle to represent basic physical laws accurately. Kang et al.[190] investigate whether large-scale video generation models can autonomously learn and generalize physical laws solely from visual data. It is found that scaling up the model size and dataset does not help with out-of-distribution physical generalization. Current models primarily rely on the existence of closely similar training examples rather than abstracting general physical rules, underscoring the need for more targeted methods to develop fully physics-aware generative models.
Cosmos Platform [26] is a recent open-source toolkit, including a video data pipeline, video tokenizers, pre-trained models, and post-trained models. Transformer-based diffusion models and transformer-based autoregressive models are trained on massive video datasets as the pre-trained world foundation models, which can be further fine-tuned on datasets for specific tasks demanding physical awareness such as robotic manipulation, camera control, and autonomous driving. Cosmos-Reason1 [191] focuses on developing multimodal large language models for embodied decisions, integrating diverse training strategies of supervised fine-tuning and reinforcement learning in terms of physical common sense, intuitive physics, and embodied reasoning. Cosmos-Reason1 also defines an ontology for physical common sense with three main categories (Space, Time, and Fundamental Physics) and 16 fine-grained categories. Cosmos-Transfer1 [192] further enables conditional world generation controlled by different spatial inputs such as segmentation, depth, and edge maps, using a ControlNet-like architecture for adaptive multimodal control.
@@
Physical Awareness from Large Language Models. Large language models can also supply physical knowledge for improving the generation of visual data. PhyT2V[177] addresses the lack of physical realism in text-to-video models by integrating large language models to refine textual prompts iteratively. The process first extracts objects and relevant physical rules from the initial text prompt, and then uses a video captioning model to generate semantic descriptions of the generated video and compare it with the original prompt to identify mismatches. Finally, the text prompt is refined based on the extracted physical properties to correct the mismatches. VideoAgent [178] produces visual plans to guide robotic systems. It refines generated video plans by leveraging self-conditioning consistency and feedback from a pretrained vision-language model. Additionally, ongoing feedback can be gathered from the environment to enable the continuous improvement of the video generation.
Physical Awareness from Physics-rich Training Data. Intentionally collected domain datasets with rich physical information can be resources for learning physical principles. WISA [20] collects the WISA-32K dataset containing approximately 32,000 videos showcasing 17 distinct physical phenomena. The dataset includes physics events across three primary domains, i.e., dynamics, thermodynamics, and optics, with detailed physical descriptions and attributes. The collected data is used to train a video generation model incorporated with a physical property embedding, a mixture-of-physical-experts attention, and a physical classifier. PISA [179] collected a dataset of real-world videos (361 videos) and synthetic videos (60 videos) showing objects dropping in diverse environments. The data is employed to improve the physical accuracy of video generation models through post-training techniques, including supervised fine-tuning and object reward optimization that encourages the alignment in terms of segmentation, optical flow, and depth maps.
@@
Generative Interactive Dynamics and Motion Controls. Generative interactive dynamics focus on modeling the natural dynamics of objects without requiring explicit physical simulations, implicitly learning the physical relationships and interactions. Blattmann et al.[180] makes an early attempt to predict object dynamics in response to localized user interactions within an image-to-video formulation. Generative Image Dynamics[181] generates realistic scene motion from a single image, via a latent diffusion model that predicts spectral volumes representing dense long-term pixel trajectories. Motion-controlled video generation and video editing belong to another form of generative dynamics[184, 59]. Motion Prompting[58] leverages motion trajectories as conditioning signals, allowing for fine-grained and flexible control of local and global scene motions. Similarly, VideoComposer[46] and Yoda[57] utilize motion vectors as explicit temporal control signals to allow for precise capture of inter-frame dynamics and patterns. Motion Dreamer[182] decouples motion reasoning from high-fidelity video generation, using intermediate motion representations such as optical flow, instance segmentation maps, and depth maps. Motion Guidance[183] controls the diffusion generation process using gradients derived from optical flow estimators, enabling detailed motion edits without model retraining.